OpenStudy (amistre64):

So I was thinking last night about 2 lines that share a common point.

OpenStudy (anonymous):

okkk

OpenStudy (amistre64):



OpenStudy (amistre64):

now, when the distance between the lines is zero, they meet and the rate of change that this happens is the difference between the slopes

OpenStudy (amistre64):

\[y_1=m_1+b_1\]\[y_2=m_2+b_2\] the distance at x=0 is b1-b2; the rate of change that effects this distance is m1-m2; therefore:\[Y_d = (m_1-m_2)X_d+(b_1-b_2)\]

OpenStudy (amistre64):

this of course is nothing new, its what happens when you apply the substitution method ....

OpenStudy (amistre64):

i just thought it was a nice way at looking at the issue :)

OpenStudy (anonymous):

Yes, it is... interesting to think about, thank you

OpenStudy (amistre64):

yw and so, the value of Xd where the distance "Yd" = 0 is then\[-\frac{b_1-b_2}{m_1-m_2}\]

OpenStudy (anonymous):

thank u amistre.. i enjoyed this short thread

OpenStudy (amistre64):

im a man of few words :)

OpenStudy (anonymous):

:)